# Constraints on the progenitor and explosion of SN 2024ggi in harmony with pre-explosion detection and hydrodynamic simulations

### v1

@@ Abstract
Supernova (SN) 2024ggi is a nearby Type II SN discovered by ATLAS, showing early flash-ionization features. The pre-explosion images reveal a red supergiant (RSG) progenitor with an initial mass of 10–17 M⊙. In the present work, we perform detailed hydrodynamic modeling to refine and put robust constraints on the progenitor and explosion parameters of SN 2024ggi. Among the progenitor models in our study, the pre-SN properties of the 11 M⊙ match the pre-explosion detected progenitor well. However, we find it difficult to completely rule out the 10 M⊙ and 12 M⊙ models. Thus, we provide a constraint of 11 M⊙ on the initial mass of the progenitor. To match the observed bolometric light curve and velocity evolution of SN 2024ggi, the favored model with an initial mass of 11 M⊙ has a pre-SN radius of 800 R⊙ and requires an explosion energy of [0.7-0.8]1051 erg, nickel mass of 0.055 M⊙, ejecta mass of 9.1 M⊙, and an amount of 0.5 M⊙ of steady-wind CSM extended up to cm resulting from an eruptive mass-loss rate of 1.0 M⊙ yr-1. We also incorporate the accelerated-wind CSM scenario which suggests a mass-loss rate of 1.0 M⊙ yr-1 and a CSM mass of 0.7 M⊙ extended up to cm. This mass-loss rate falls within the range constrained observationally. Additionally, due to the constraint of 11 M⊙ on the initial mass, the range of pre-SN radius and ejecta mass would be [690–900] R⊙, and [8.2–9.6] M⊙, respectively.
@@ 1 Introduction
Followed by the discovery of the core-collapse supernova (CCSN) 2023ixf (K. Itagaki, 2023) at a distance less than 10 Mpc (DLT10), another DLT10 CCSN 2024ggi was discovered in NGC 3621 by the Asteroid Terrestrial-impact Last Alert System (ATLAS) on the 11th of April 2024 (JD=2460411.64069, J. Tonry et al., 2024; S. Srivastav et al., 2024; T.-W. Chen et al., 2025) and it was classified as Type II (W. Hoogendam et al., 2024; Q. Zhai et al., 2024). Immediate spectroscopic follow-up starting only a few hours since discovery revealed the photo-ionized, optically-thick CSM surrounding the exploding star through the presence of strong, flash-ionization features in the spectra (W. V. Jacobson-Galán et al., 2024; T. Pessi et al., 2024; M. Shrestha et al., 2024; J. Zhang et al., 2024; T.-W. Chen et al., 2025).
Due to its proximity, the site of SN 2024ggi had been visited by numerous telescopes on several occasions before the explosion. Utilizing the pre-explosion images from the Hubble Space Telescope (HST) and the Spitzer Space Telescope, D. Xiang et al. (2024) identified a red, bright variable star with varying absolute -band magnitudes of mag in 1995 to mag in 2003. The variation in -band magnitudes was found consistent with a normal RSG star. For the identified progenitor, D. Xiang et al. (2024) constrained the temperature and bolometric luminosity to be = K and log(/L⊙) = , respectively.
In another study, X. Chen et al. (2024) examined the multi-band deep archival images in -bands from the Dark Energy Camera Legacy Survey, and provided the range of progenitor mass lying in 14–17 M⊙. However, their progenitor mass estimation could be influenced by the flux contamination from a nearby source, as S. Srivastav et al. (2024) had indicated that the red source was resolved into two sources in -band image from HST. S. Yang et al. (2024) also measured the brightness at the location of SN 2024ggi in the DESI Legacy Survey images and found absolute magnitudes within the range typical for an RSG. Further utilizing the HST images, X. Hong et al. (2024) put constraints on the progenitor mass through independent measurements unaffected by the circumstellar extinction and pulsational brightness variability. They found the dying star associated with the youngest population in SN environment with an age (log[/yr]) of 7.41, corresponding to an initial mass of 10.2 M⊙.
Despite some constraints over the progenitor masses with the direct detection of the progenitor in pre-explosion images, detailed hydrodynamic simulations are lacking in the literature to strengthen the claims on the progenitor star and explosion properties. However, K. Ertini et al. (2025) utilized a grid of progenitor models available from K. Nomoto & M. Hashimoto (1988) to reproduce the observed bolometric light curve and velocity evolution of SN 2024ggi. They constrained an initial progenitor mass of 15 M⊙.
@@ 2 Constraints on the possible progenitor
As seen above, initially, the direct detection of the possible progenitor in the pre-explosion images put a constraint of 13 – 17 M⊙ (D. Xiang et al., 2024; X. Chen et al., 2024) on the progenitor mass. However, an environmental analysis (free from the effect of circumstellar extinction and pulsational brightness variability) by X. Hong et al. (2024) utilizing the pre-explosion HST images of SN 2024ggi site, places the progenitor towards the slightly lower mass regime. As a result, we choose [10–17] M⊙, ZAMS star models in our simulations to serve as the possible progenitor for SN 2024ggi.
@@ 2.1 Details on simulation setups to create pre-SN models
Starting from the pre-main-sequence, our models arrive on the ZAMS with specified masses and respective rotations, pass through stages of successive nuclear burning and finally reach the pre-SN stages marked by the Iron-core (Fe-core) infall. To evolve the models upto the stages of Fe-core infall, we employ the state-of-the-art simulation tool MESA, version r24.03.1 (B. Paxton et al., 2011, 2013, 2015, 2018, 2019; A. S. Jermyn et al., 2023). Our models have solar metallicity (Z = 0.0142, following M. Asplund et al. 2009) and nominal rotation, which is only the 5% of their critical angular rotational velocity. The solar metallicity composition for SN 2024ggi progenitor is supported by W. V. Jacobson-Galán et al. (2024); D. Xiang et al. (2024); X. Chen et al. (2024); T.-W. Chen et al. (2025); X. Hong et al. (2024). The nominal angular rotation velocity of only 5% of their critical angular rotational velocity is supported by the fact that the progenitors of Type IIP SNe are RSGs (confirmed through direct detection of progenitors, e.g., S. J. Smartt 2009; S. D. Van Dyk 2017) that are typically very slow or non rotating due to their extended envelope (A. Heger & N. Langer, 1998; A. Maeder & G. Meynet, 2001).
@@
We use the 12M_pre_ms_to_core_collapse test-suite directory of MESA to create the pre-SN models. As specified in the recent versions of MESA setups, when the hydrogen mass fraction in the core decreases to 0.68, the model is assumed to have reached the ZAMS stage (previously, the ratio of energy produced from nuclear reactions and the overall luminosity of the model being 0.4–0.8 was used to define the arrival on ZAMS). In our models, convection is modeled following the mixing theory of L. Henyey et al. (1965), adopting the Ledoux criterion. We have adopted a mixing-length-theory parameter () = 2.0 (typical value being 1–3, as used in R. Farmer et al., 2016; A. Aryan et al., 2021; R. Ouchi & K. Maeda, 2021; S. B. Pandey et al., 2021; A. Aryan et al., 2022c, b, a, 2023; A. K. Ror et al., 2024). Following N. Langer et al. (1985), semi-convection is incorporated in our models with an efficiency parameter of = 0.01. To model the convective overshooting, we use the default MESA settings for massive stars by setting overshooting parameters, and for the hydrogen core, assuming step overshooting. The choice of step overshooting for the hydrogen core in massive stars introduces a sharp boundary extension that ultimately helps capture the increased main-sequence lifetime and observed main-sequence width. The chosen value of and using step overshooting for hydrogen core are supported by I. Brott et al. (2011) and also by asteroseismology results around this mass range (D. M. Bowman, 2020). The helium core would be relatively smaller, less vigorous, and the overshoot-induced mixing is expected to decay with distance. Thus, we use and for helium core assuming exponential overshooting utilizing the results of F. Herwig (2000). In addition to these convective core overshooting, we also use a small amount of overshooting on the top of any other convective core to avoid spurious numerical behavior. The RSGs retain most of their outer envelope; thus, we use the ‘Dutch’ wind scheme (E. Glebbeek et al., 2009) with scaling factor () of 0.1 only, which accounts for the small mass losses. This wind scheme incorporates the outcomes from multiple works for different situations; (a) in the regions with the effective temperature, K along with hydrogen mass fraction 0.4, the outcomes of J. S. Vink et al. (2001) are used; (b) when K and hydrogen mass fraction 0.4, the results of T. Nugis & H. J. G. L. M. Lamers (2000) are used; and (c) the wind scheme presented in C. de Jager et al. (1988) is used for the regions with K. There are several instances where produced the desired amount of hydrogen mass in pre-explosion RSG models for Type IIP SNe (e.g., for SN 2012A, B. Paxton et al. 2018 used assuming Z = 0.02; also for SN 2005cs but with Z = 0.006). These studies of SN 2012A and SN 2005cs have motivated our choice of wind efficiency. Throughout the simulation, we use appropriate temporal and spatial resolution. Figure 1 displays the effect of changing the temporal and spatial resolutions for the 11 M⊙, thus confirming the convergence of our model grid within the assumed resolutions in our study.
@@ 2.2 Discussion on the choice of and rotation
The progenitors of Type IIP SNe are RSGs with their outer hydrogen envelope almost intact. We have assumed a small to account for the low mass-loss rates. There are examples (e.g., B. Paxton et al., 2018) where an has helped to produce the desired amounts of hydrogen and ejecta masses to explain the observed plateau durations of Type IIP SNe. In Figure 2, we see that changing from 0.1 to 0.2 for the 15 M⊙ has a negligible effect on the overall evolution of the star. Their physical properties at different stages of evolution are also almost identical, as presented in Table 1. Additionally, it is argued that RSG mass-loss rates may be much lower than the ‘Dutch’ prescription (E. R. Beasor et al., 2020). Further, while modeling the light curves of SN 2004A, SN 2004et, SN 2009ib, SN 2017eaw, and SN 2017gmr, J. A. Goldberg & L. Bildsten (2020) found several models with explaining the bolometric light curves and velocity evolution nicely. We also investigate the effect of changing to 1. One important thing to notice in Figure 2 is that the terminating positions of the models with different on the HR-diagram do not change considerably, as shown by the 11 M⊙ and 17 M⊙ models. However, due to an increased , the pre-SN mass changes considerably as presented in Table 1.
@@
Rotation in stars is a natural consequence of angular momentum conservation during the gravitational collapse of a star-forming molecular cloud. Virtually all stars rotate, although the rotation rate varies widely depending on physical characteristics like mass, age, metallicity, magnetic braking, and evolutionary stage. Several studies have shown that including rotation in stellar evolution models helps to explain the observed surface chemical enrichment, extended lifetimes, position on HR-diagram, and pre-SN core masses (among several others, A. Maeder & G. Meynet, 2000; I. Hunter et al., 2008; C. Georgy et al., 2013). As indicated by S. Ekström et al. (2012), for the stars with initial mass range in 9–20 M⊙, the surface rotation velocities after the end of their respective core carbon burning stages are virtually close to zero. In the present study, we choose an initial angular rotational velocity of only 5% of the critical angular rotational velocity. As expected, the final surface velocities in our models are exceptionally low ( 0.02 km s-1 for the 11 M⊙). We also show the non-rotating model for the 11 M⊙ and find that its evolution is almost identical to the slow rotating case (i.e., 5% Rot case) within the limits of the considered temporal and spatial resolutions here. One can argue for choosing high initial angular rotational velocities, however, as shown for the 14 M⊙ in Figure 2, increasing the initial angular rotational velocity from 5% to 40% significantly alters the overall evolution of the models. A critical consequence of increasing the initial rotation would be that the terminating stages of the models with high initial masses would move further away on the HR-diagram from the pre-explosion detected progenitor. Thus, a rapidly rotating high-mass progenitor would be less favored for SN 2024ggi.
@@
Additional models: 11 M⊙ with fixed =0.055 M⊙ and different | |||||||||||
1.73 | 9.1 | 0.055 | 0.6 | ||||||||
1.69 | 9.1 | 0.055 | 0.8 | ||||||||
@@
1.71 | 9.1 | 0.035 | 0.7 | ||||||||
1.71 | 9.1 | 0.045 | 0.7 | ||||||||
1.71 | 9.1 | 0.065 | 0.7 |
a Mass at ZAMS. bProgenitor radius at ZAMS, cLuminosity at ZAMS, dFinal mass at the pre-SN stage, ePre-SN phase radius, fPre-SN phase luminosity, gFinal core mass after fallback, hEjecta mass, iAmount of synthesized Nickel used in the explosion, jtotal energy after explosion energy. † The non-rotating case. The properties of the non-rotating 11 M⊙ model are indistinguishable with the slowly rotating model within the limits of resolutions considered here. It’s explosion results are expected to be almost similar as the slow rotating case (i.e., similar to Rot 5% case). ∗ With = 0.2. ∗∗ With = 1. ‡ With initial angular rotation velocity of 40 of critical angular rotation velocity. Unless mentioned, = 0.1 is used. We plan to put all the MESA-inlist-files used to generate our models on Zenodo.

### v2

@@ Abstract
Supernova (SN) 2024ggi is a nearby Type II SN discovered by ATLAS, showing early flash-ionization features. The pre-explosion images reveal a red supergiant (RSG) progenitor with an initial mass of 10–17 M⊙. In the present work, we perform detailed hydrodynamic modeling to refine and put robust constraints on the progenitor and explosion parameters of SN 2024ggi. Among the progenitor models in our study, the pre-SN properties of the 11 M⊙ match the pre-explosion detected progenitor well. However, we find it difficult to completely rule out the 10 M⊙ and 12 M⊙ models. Thus, we provide a constraint of 11 M⊙ on the initial mass of the progenitor. To match the observed bolometric light curve and velocity evolution of SN 2024ggi, the favored model with an initial mass of 11 M⊙ has a pre-SN radius of 800 R⊙. This model requires an explosion energy of [0.7-0.8]1051 erg, nickel mass of 0.049 M⊙, ejecta mass of 9.1 M⊙, and an amount of 0.5 M⊙ of steady-wind CSM extended up to cm resulting from an eruptive mass-loss rate of 1.0 M⊙ yr-1. We also incorporate the accelerated-wind CSM scenario which suggests a mass-loss rate of 1.0 M⊙ yr-1 and a CSM mass of 0.7 M⊙ extended up to cm. This mass-loss rate falls within the range constrained observationally. Additionally, due to the constraint of 11 M⊙ on the initial mass, the range of pre-SN radius and ejecta mass would be [690–900] R⊙, and [8.2–9.6] M⊙, respectively.
@@ 1 Introduction
Followed by the discovery of the core-collapse supernova (CCSN) 2023ixf (2023TNSTR1158....1I) at a distance less than 10 Mpc (DLT10), another DLT10 CCSN 2024ggi was discovered in NGC 3621 by the Asteroid Terrestrial-impact Last Alert System (ATLAS) on the 11th of April 2024 (JD=2460411.64069, 2024TNSTR1020....1T; 2024TNSAN.100....1S; 2025ApJ...983...86C) and it was classified as Type II (2024TNSAN.103....1H; 2024TNSCR1031....1Z). Immediate spectroscopic follow-up starting only a few hours since discovery revealed the photo-ionized, optically-thick CSM surrounding the exploding star through the presence of strong, flash-ionization features in the spectra (2024ApJ...972..177J; 2024A&A...688L..28P; 2024ApJ...972L..15S; 2024ApJ...970L..18Z; 2025ApJ...983...86C).
Due to its proximity, the site of SN 2024ggi had been visited by numerous telescopes on several occasions before the explosion. Utilizing the pre-explosion images from the Hubble Space Telescope (HST) and the Spitzer Space Telescope, 2024ApJ...969L..15X identified a red, bright variable star with varying absolute -band magnitudes of mag in 1995 to mag in 2003. The variation in -band magnitudes was found consistent with a normal RSG star. For the identified progenitor, 2024ApJ...969L..15X constrained the temperature and bolometric luminosity to be = K and log(/L⊙) = , respectively.
In another study, 2024ApJ...971L...2C examined the multi-band deep archival images in -bands from the Dark Energy Camera Legacy Survey, and provided the range of progenitor mass lying in 14–17 M⊙. However, their progenitor mass estimation could be influenced by the flux contamination from a nearby source, as 2024TNSAN.100....1S had indicated that the red source was resolved into two sources in -band image from HST. 2024TNSAN.105....1Y also measured the brightness at the location of SN 2024ggi in the DESI Legacy Survey images and found absolute magnitudes within the range typical for an RSG. Further utilizing the HST images, 2024ApJ...977L..50H put constraints on the progenitor mass through independent measurements unaffected by the circumstellar extinction and pulsational brightness variability. They found the dying star associated with the youngest population in SN environment with an age (log[/yr]) of 7.41, corresponding to an initial mass of 10.2 M⊙.
Despite some constraints over the progenitor masses through the direct detection of the possible progenitor in pre-explosion images, detailed hydrodynamic simulations are scarce in the literature to strengthen the claims on the progenitor star and explosion properties. However, 2025arXiv250301577E utilized a grid of progenitor models available from 1988PhR...163...13N to reproduce the observed bolometric light curve and velocity evolution of SN 2024ggi. They constrained an initial progenitor mass of 15 M⊙.
@@ 2 Constraints on the possible progenitor
As seen above, initially, the direct detection of the possible progenitor in the pre-explosion images put a constraint of 13 – 17 M⊙ (2024ApJ...969L..15X; 2024ApJ...971L...2C) on the progenitor mass. However, an environmental analysis (free from the effect of circumstellar extinction and pulsational brightness variability) by 2024ApJ...977L..50H utilizing the pre-explosion HST images of SN 2024ggi site, places the progenitor towards the slightly lower mass regime. As a result, we choose [10–17] M⊙, ZAMS star models in our simulations to serve as the possible progenitor for SN 2024ggi.
@@ 2.1 Details on simulation setups to create pre-SN models
Starting from the pre-main-sequence, our models arrive on the ZAMS with specified masses and respective rotations, pass through stages of successive nuclear burning and finally reach the pre-SN stages marked by the Iron-core (Fe-core) infall. To evolve the models upto the stages of Fe-core infall, we employ the state-of-the-art simulation tool MESA, version r24.03.1 (2011ApJS..192....3P; 2013ApJS..208....4P; 2015ApJS..220...15P; 2018ApJS..234...34P; 2019ApJS..243...10P; 2023ApJS..265...15J). Our models have solar metallicity (Z = 0.0142, following 2009ARA&A..47..481A) and nominal rotation, which is only the 5% of their critical angular rotational velocity. The solar metallicity composition for SN 2024ggi progenitor is supported by 2024ApJ...972..177J; 2024ApJ...969L..15X; 2024ApJ...971L...2C; 2025ApJ...983...86C; 2024ApJ...977L..50H. The nominal angular rotation velocity of only 5% of their critical angular rotational velocity is supported by the fact that the progenitors of Type IIP SNe are RSGs (confirmed through direct detection of progenitors, e.g., 2009ARA&A..47...63S; VanDyk2017) that are typically very slow or non rotating due to their extended envelope (1998A&A...334..210H; 2001A&A...373..555M).
@@
We use the 12M_pre_ms_to_core_collapse test-suite directory of MESA to create the pre-SN models. As specified in the recent versions of MESA setups, when the hydrogen mass fraction in the core decreases to 0.68, the model is assumed to have reached the ZAMS stage (previously, the ratio of energy produced from nuclear reactions and the overall luminosity of the model being 0.4–0.8 was used to define the arrival on ZAMS). In our models, convection is modeled following the mixing theory of 1965ApJ...142..841H, adopting the Ledoux criterion. We have adopted a mixing-length-theory parameter () = 2.0 (typical value being 1–3, as used in 2016ApJS..227...22F; 2021MNRAS.505.2530A; 2021MNRAS.500.1889O; 2021MNRAS.507.1229P; 2022MNRAS.517.1750A; 2022JApA...43....2A; 2022JApA...43...87A; 2023MNRAS.521L..17A; 2024ApJ...971..163R). Following 1985A&A...145..179L, semi-convection is incorporated in our models with an efficiency parameter of = 0.01. To model the convective overshooting, we use the default MESA settings for massive stars by setting overshooting parameters, and for the hydrogen core, assuming step overshooting. The choice of step overshooting for the hydrogen core in massive stars introduces a sharp boundary extension that ultimately helps capture the increased main-sequence lifetime and observed main-sequence width. The chosen value of and using step overshooting for hydrogen core are supported by 2011A&A...530A.115B and also by asteroseismology results around this mass range (2020FrASS...7...70B). The helium core would be relatively smaller, less vigorous, and the overshoot-induced mixing is expected to decay with distance. Thus, we use and for helium core assuming exponential overshooting utilizing the results of 2000A&A...360..952H. In addition to these convective core overshooting, we also use a small amount of overshooting on the top of any other convective core to avoid spurious numerical behavior. The RSGs retain most of their outer envelope; thus, we use the ‘Dutch’ wind scheme (2009A&A...497..255G) with scaling factor () of 0.1 only, which accounts for the small mass losses. This wind scheme incorporates the outcomes from multiple works for different situations; (a) in the regions with the effective temperature, K along with hydrogen mass fraction 0.4, the outcomes of 2001A&A...369..574V are used; (b) when K and hydrogen mass fraction 0.4, the results of 2000A&A...360..227N are used; and (c) the wind scheme presented in 1988A&AS...72..259D is used for the regions with K. Throughout the simulation, we use appropriate temporal and spatial resolution. Figure 1 displays the effect of changing the temporal and spatial resolutions for the 11 M⊙, thus confirming the convergence of our model grid within the assumed resolutions in our study.
@@ 2.2 Discussion on the choice of and rotation
The progenitors of Type IIP SNe are RSGs with their outer hydrogen envelope almost intact. We have assumed a small to account for the low mass-loss rates. There are examples (e.g., 2018ApJS..234...34P) where an has helped to produce the desired amounts of hydrogen and ejecta masses to explain the observed plateau durations of Type IIP SNe. In Figure 2, we see that changing from 0.1 to 0.2 for the 11 M⊙ (and also 15 M⊙) has a negligible effect on the overall evolution of the star. Their physical properties at different stages of evolution are also almost identical, as presented in Table 1. Additionally, it is argued that RSG mass-loss rates may be much lower than the ‘Dutch’ prescription (2020MNRAS.492.5994B). We also refer to 2025Galax..13...72V for a review on RSG mass losses, which supports the typically modest rates of mass loss for progenitors of SNe IIP, compared to those of more massive red supergiants. Further, while modeling the light curves of SN 2004A, SN 2004et, SN 2009ib, SN 2017eaw, and SN 2017gmr, 2020ApJ...895L..45G found several models with explaining the bolometric light curves and velocity evolution nicely. We also investigate the effect of changing to 1. One important thing to notice in Figure 2 is that the terminating positions of the models with different on the HR-diagram do not change considerably, as shown by the 11 M⊙ and 17 M⊙ models. However, due to an increased , the pre-SN mass changes considerably as presented in Table 1.
@@
Rotation in stars is a natural consequence of angular momentum conservation during the gravitational collapse of a star-forming molecular cloud. Virtually all stars rotate, although the rotation rate varies widely depending on physical characteristics like mass, age, metallicity, magnetic braking, and evolutionary stage. Several studies have shown that including rotation in stellar evolution models helps to explain the observed surface chemical enrichment, extended lifetimes, position on HR-diagram, and pre-SN core masses (among several others, 2000ARA&A..38..143M; 2008A&A...479..541H; 2013A&A...558A.103G). As indicated by 2012A&A...537A.146E, for the stars with initial mass range in 9–20 M⊙, the surface rotation velocities after the end of their respective core carbon burning stages are virtually close to zero. In the present study, we choose an initial angular rotational velocity of only 5% of the critical angular rotational velocity. As expected, the final surface velocities in our models are exceptionally low ( 0.02 km s-1 for the 11 M⊙). We also show the non-rotating model for the 11 M⊙ and find that its evolution is almost identical to the slow rotating case (i.e., 5% Rot case) within the limits of the considered temporal and spatial resolutions here. One can argue for choosing high initial angular rotational velocities, however, as shown for the 14 M⊙ in Figure 2, increasing the initial angular rotational velocity from 5% to 40% significantly alters the overall evolution of the models. A critical consequence of increasing the initial rotation would be that the terminating stages of the models with high initial masses would move further away on the HR-diagram from the pre-explosion detected progenitor. Thus, a rapidly rotating high-mass progenitor would be less favored for SN 2024ggi.
@@
| Additional models: 11 M⊙ with fixed =0.049 M⊙ and different | |||||||||||
| 1.73 | 9.1 | 0.049 | 0.6 | ||||||||
| 1.69 | 9.1 | 0.049 | 0.8 | ||||||||
@@
| 1.71 | 9.1 | 0.035 | 0.7 | ||||||||
| 1.71 | 9.1 | 0.045 | 0.7 | ||||||||
| 1.71 | 9.1 | 0.049 | 0.7 | ||||||||
| 1.71 | 9.1 | 0.055 | 0.7 | ||||||||
| 1.71 | 9.1 | 0.065 | 0.7 |
a Mass at ZAMS. bProgenitor radius at ZAMS, cLuminosity at ZAMS, dFinal mass at the pre-SN stage, ePre-SN phase radius, fPre-SN phase luminosity, gFinal core mass after fallback, hEjecta mass, iAmount of synthesized Nickel, jtotal energy after explosion energy. † The non-rotating case. The properties of the non-rotating 11 M⊙ model are indistinguishable with the slowly rotating model within the limits of resolutions considered here. It’s explosion results are expected to be almost similar as the slow rotating case (i.e., similar to Rot 5% case). ∗ With = 0.2. ∗∗ With = 1. ‡ With initial angular rotation velocity of 40 of critical angular rotation velocity. Unless mentioned, = 0.1 is used. All the MESA-inlist-files, pre-SN models, STELLA input files to generate luminosity, velocity evolution, and also the final STELLA result files, are available on https://zenodo.org/records/17396113 (catalog Zenodo, MESA community) and also on https://zenodo.org/records/17395362 (catalog Zenodo, AAS Journals community).